### 18 Mar 2026 v1

# Improving causal inference in interrupted time series analysis: the triple difference design

Abstract
Background: Interrupted time series analysis (ITSA) is a valuable quasi-experimental design for evaluating health policy and intervention effects. While multiple-group ITSA (MG-ITSA) strengthens causal inference by incorporating a control group, residual confounding from unmeasured time-varying factors may persist. The triple-difference interrupted time series (DDD-ITSA) design extends this framework by adding a second control group to further isolate treatment effects, yet the design remains underutilized and lacks formal methodological guidance.
Methods: This paper formalizes the DDD-ITSA design for health research. We specify the underlying regression model, define the key coefficients for estimating level and trend effects, and clarify the interpretation of the triple-difference estimand. We also illustrate application of the design using a worked example evaluating the effect of California’s Proposition 99 cigarette tax increase on per-capita cigarette sales.
Results: The DDD-ITSA model provides clear parameterization for assessing baseline balance across groups and estimating differential treatment effects. In the stylized Proposition 99 example, all groups demonstrated balance on pre-intervention level and trend. The triple-difference estimand revealed a statistically significant annual net reduction of -1.76 per-capita cigarette packs in California relative to the secondary control (P = 0.020; 95% CI: -3.24, -0.28), consistent with the treatment effect observed against the primary control. The difference-in-differences between the two control groups was non-significant, supporting the validity of the comparison.
Conclusions: The DDD-ITSA design offers a robust approach for strengthening causal inferences when two-group comparisons may be confounded. By leveraging an additional control group, researchers can difference out remaining biases and assess effect heterogeneity across subgroups. The updated itsa package for Stata facilitates implementation, making this design accessible for routine application in health policy research. Careful attention to control group selection, baseline balance, and autocorrelation remains essential for valid interpretation.
Keywords
difference-in-difference-in-differences, triple differences, interrupted time series analysis, controlled interrupted time series analysis, multiple-group interrupted time series analysis
1 Background
Interrupted time series analysis (ITSA) is widely used in healthcare research to estimate the causal effects of interventions, policy changes, and natural events when randomized designs are infeasible. The approach relies on sequential outcome measurements collected at equally spaced intervals, typically aggregated at the unit level (e.g., hospitals, regions, or populations) and summarized as rates, counts, or mean expenditures. In its canonical form, a single unit is exposed to an intervention that is hypothesized to generate a structural change (“interruption”) in the level and/or slope of the outcome process at a known point in time [4, 24]. Identification in this framework is achieved by using the pre-intervention segment of the series to extrapolate the counterfactual post-intervention trajectory under the assumption that, absent the intervention, the underlying data-generating process would have continued unchanged. The use of multiple pre-intervention observations partially addresses threats such as regression to the mean and improves internal validity relative to simpler before–after comparisons [4, 24, 13].
Despite these strengths, single-group ITSA (SG-ITSA) rests on the strong identifying assumption that no unmeasured, time-varying confounders or coincident events affect the outcome at the time of intervention. When this assumption is violated (such as in the presence of secular trends, anticipatory behavior, or concurrent policy changes), SG-ITSA may produce biased estimates of the intervention effect. Empirical evidence indicates that SG-ITSA can yield misleading inferences when pre-intervention trends already exhibit directional change or when external influences evolve contemporaneously with the intervention [12, 16, 17].
Multiple-group or controlled ITSA (MG-ITSA) addresses these threats by introducing control units that were not exposed to the intervention but share similar baseline levels, pre-intervention trends, and observed characteristics with the treated unit [14, 1, 18]. By differencing outcomes between treated and control series over time, MG-ITSA relaxes the strong assumption of stable counterfactual trends within a single unit and instead relies on a parallel-trends-type assumption across groups [14, 18]. Under this framework, causal effects are identified by removing secular trends and common shocks that affect all units similarly, aligning MG-ITSA conceptually with difference-in-differences estimators [23]. However, when treatment effects are expected to vary across an additional dimension, such as subpopulations, exposure intensity, or policy eligibility, standard MG-ITSA may remain insufficient to isolate the causal effect of interest.
A triple-difference interrupted time series (DDD-ITSA) design extends the logic of the MG-ITSA by adding a second control group for comparison. This approach mirrors the difference-in-differences-in-differences (DDD) framework in the econometrics literature, where an additional difference is used to remove remaining confounding that may persist after a standard two-group comparison [21]. In the ITSA context, the triple-difference estimator captures whether the post-intervention change in level and/or trend for the treated group is larger (or smaller) for the subgroup of interest than would be expected based on concurrent changes in control groups and comparison subgroups. By differencing across time, groups, and subgroups simultaneously, the DDD-ITSA design strengthens causal identification under assumptions analogous to those in the difference-in-differences (DiD) literature, namely that, absent the intervention, these differences would have evolved similarly over time [23].
Applications that explicitly implement a DDD-ITSA framework remain relatively uncommon in the literature. One applied example is the evaluation of Maryland’s global budget revenue (GBR) model, where researchers used an interrupted time-series design with difference-in-differences comparisons and a triple differences analysis to assess changes in emergency department utilization, admissions, and revisits following GBR implementation, including how these trends varied across racial/ethnic and payer groups [7]. In this study, the triple-difference component helped identify differential effects of the policy on emergency department return outcomes across subpopulations beyond general statewide trends and group-level effects. Another example is research on the effects of the Affordable Care Act (ACA) on health insurance coverage among Latino subgroups, which combined ITSA with triple-difference models to evaluate heterogeneous changes in uninsured rates associated with the ACA and Medicaid expansion across multiple Latino ethnic groups [8].
This paper formalizes the DDD-ITSA framework for applied health research. We describe the design, detail the underlying statistical model, clarify the interpretation of the key coefficients that identify level and trend effects, and present a worked empirical example to illustrate implementation and interpretation. We also provide practical guidance regarding identification assumptions, model specification, and appropriate use cases for the design. To facilitate applied use, the itsa package for Stata [14] has been updated to estimate DDD-ITSA models, enabling researchers to easily implement the approach using their data.
2 Methods
2.1 The DDD-ITSA design
The DDD-ITSA design is an extension of the MG-ITSA design to an additional control group that is either geographically distant from the treatment and primary control groups, or comprises a different demographic that still allows for comparison (e.g., type of insurance coverage, different medical condition, ethnicity, etc.). This second control group may be thought of as adding a “doubly robustness” property to the analysis, in that it provides a second opportunity to get the causal inference correct if the treatment versus primary control is confounded. Figure 1 illustrates four of the many potential outcome scenarios that we could anticipate in the DDD-ITSA design when emphasizing differences in trends: (A) a treatment effect in the primary analysis and secondary analysis; (B) no treatment effect in either the primary or secondary analysis; (C) a treatment effect in the secondary analysis but not in the primary analysis (evidence of confounding); and (D) a treatment effect in the primary analysis but not in the secondary analysis (suggesting external influences or spillover affecting the secondary control).
Figure 1: Four general outcome scenarios for the DDD-ITSA (A) treatment effect in both primary and secondary analyses, (B) no treatment effect, (C) treatment effect in secondary analysis only, (D) treatment effect in primary only.
2.2 The DDD-ITSA model
The DDD-ITSA regression model is an extension of the MG-ITSA [14, 15, 16] and assumes the following form:
$Y_{t}$ $=\beta_{0}+\beta_{1}T_{t}+\beta_{2}X_{t}+\beta_{3}X_{t}T_{t}+\beta_{4}Z_{1}+\beta_{5}Z_{1}T_{t}$ (1)
$\quad+\beta_{6}Z_{1}X_{t}+\beta_{7}Z_{1}X_{t}T_{t}+\beta_{8}Z_{2}+\beta_{9}Z_{2}T_{t}$
$\quad+\beta_{10}Z_{2}X_{t}+\beta_{11}Z_{2}X_{t}T_{t}+\epsilon_{t}$
where $Y_{t}$ is the aggregated outcome variable measured at each time-point $t$, $T_{t}$ is the time since the start of the study, $X_{t}$ is a dummy variable representing the treatment (pre-treatment periods = 0, otherwise 1), $Z_{1}$ is a dummy variable to denote the treatment unit and $Z_{2}$ is a dummy variable to denote the second control group. Therefore, all variables and coefficients that include $Z_{1}$ are equivalent to those in an MG-ITSA model in which the treatment unit is compared to the control group, and all variables and coefficients that include $Z_{2}$ compare the second control group to the first control group.
More specifically, $\beta_{4}$ represents the difference in the level (intercept) of the dependent variable between treatment and the first control group prior to the intervention, $\beta_{5}$ represents the difference in the slope (trend) of the dependent variable between treatment and the first control group prior to the intervention, $\beta_{6}$ indicates the difference between treatment and the first control group in the level of the dependent variable immediately following introduction of the intervention, and $\beta_{7}$ represents the difference between treatment and the first control group in the slope (trend) of the dependent variable after initiation of the intervention compared with pre-intervention.
Similarly, $\beta_{8}$ represents the difference in the level (intercept) of the dependent variable between the second control group and the first control group prior to the intervention, $\beta_{9}$ represents the difference in the slope (trend) of the dependent variable between the second control and the first control group prior to the intervention, $\beta_{10}$ indicates the difference between the second control group and the first control group in the level of the dependent variable immediately following introduction of the intervention, and $\beta_{11}$ represents the difference between the second control group and the first control group in the slope (trend) of the dependent variable after initiation of the intervention compared with pre-intervention.
When the random error terms follow an AR(1) process,
$\epsilon_{t}=\rho\epsilon_{t-1}+u_{t}$ (2)
where the autocorrelation parameter $\rho$ is the correlation coefficient between adjacent error terms such that $|\rho|<1$, and the disturbances $u_{t}$ are independent $N(0,\sigma^{2})$ [10].
Table 1 presents the coefficients and linear combination terms for the pre-treatment and post-treatment trends of each of the three groups, and Table 2 presents the difference-in-differences of trend comparisons between groups.
Table 1: Pre-treatment and post-treatment trends, by group
Group Pre-treatment (X = 0) Post-treatment (X = 1) Pre–post trend change
Control 1 (Z1 = 0, Z2 = 0) $\beta_{1}$ $\beta_{1}+\beta_{3}$ $\beta_{3}$
Treatment (Z1 = 1, Z2 = 0) $\beta_{1}+\beta_{5}$ $\beta_{1}+\beta_{3}+\beta_{5}+\beta_{7}$ $\beta_{3}+\beta_{7}$
Control 2 (Z2 = 1, Z1 = 0) $\beta_{1}+\beta_{9}$ $\beta_{1}+\beta_{3}+\beta_{9}+\beta_{11}$ $\beta_{3}+\beta_{11}$
Table 2: Difference-in-differences (trend change) comparisons between groups
Comparison DiD Simplified expression
Treatment vs Control 1 $(\beta_{3}+\beta_{7})-\beta_{3}$ $\beta_{7}$
Control 2 vs Control 1 $(\beta_{3}+\beta_{11})-\beta_{3}$ $\beta_{11}$
Treatment vs Control 2 $(\beta_{3}+\beta_{7})-(\beta_{3}+\beta_{11})$ $\beta_{7}-\beta_{11}$
Tables 3 and 4 present the analogous coefficients and linear combination terms for the pre-treatment and post-treatment level changes of each of the three groups and their respective difference in differences.
Table 3: Pre-treatment and post-treatment levels, by group
Group Pre-treatment (X = 0) Post-treatment (X = 1) Pre–post level change
Control 1 (Z1 = 0, Z2 = 0) $\beta_{0}$ $\beta_{0}+\beta_{2}$ $\beta_{2}$
Treatment (Z1 = 1, Z2 = 0) $\beta_{0}+\beta_{4}$ $\beta_{0}+\beta_{2}+\beta_{4}+\beta_{6}$ $\beta_{2}+\beta_{6}$
Control 2 (Z2 = 1, Z1 = 0) $\beta_{0}+\beta_{8}$ $\beta_{0}+\beta_{2}+\beta_{8}+\beta_{10}$ $\beta_{2}+\beta_{10}$
Table 4: Difference-in-differences (level change) comparisons between groups
Comparison DiD Simplified expression
Treatment vs Control 1 $(\beta_{2}+\beta_{6})-\beta_{2}$ $\beta_{6}$
Control 2 vs Control 1 $(\beta_{2}+\beta_{10})-\beta_{2}$ $\beta_{10}$
Treatment vs Control 2 $(\beta_{2}+\beta_{6})-(\beta_{2}+\beta_{10})$ $\beta_{6}-\beta_{10}$
2.3 Assessing balance on baseline level and trend
In non-experimental settings, it is critical to evaluate whether the treatment and control groups are balanced in both the level and slope of the outcome variable prior to the intervention. Balance across groups in baseline levels and trends can be formally assessed using the coefficients from Equation (1). Specifically, statistically insignificant estimates of $\beta_{4}$ and $\beta_{5}$ indicate no systematic differences between the treatment group and the primary control group in pre-intervention levels and trends, respectively. Likewise, statistical insignificance of the linear combinations $(\beta_{4}-\beta_{8})$ and $(\beta_{5}-\beta_{9})$ suggests baseline equivalence between the treatment group and the secondary control group in levels and trends. Finally, statistically insignificant $\beta_{8}$ and $\beta_{9}$ imply balance between the two control groups in pre-treatment levels and trajectories.
Although multiple baseline trend comparisons are evaluated, Olden and Møen [21] demonstrate that identification of the DDD estimator does not require two independent parallel trends assumptions. The intuition is that the DDD estimator can be interpreted as the difference between two potentially biased difference-in-differences estimators. As long as the bias component is identical across these estimators, it differences out, yielding an unbiased DDD estimate. Consequently, only a single parallel trends condition is required for causal interpretation.
2.4 Example
The following example illustrates the analysis of the DDD-ITSA design using data from California’s 1988 enactment of Proposition 99, a voter-approved initiative that increased the cigarette excise tax by \$0.25 per pack and funded statewide anti-smoking efforts. The outcome examined is per-capita cigarette sales (packs), measured annually at the state level from 1970 to 2000, with 1989 marking the first post-intervention year. The dataset includes cigarette sales and four covariates: average retail cigarette price, logged per-capita personal income, per-capita beer consumption, and the share of the population aged 15–24.
For exposition purposes, Idaho and Montana are assigned to the primary control group and Colorado as the secondary control state. These states were identified as balanced matches to California (the treated state) on baseline level and trend of cigarette sales using the itsamatch program in Stata [1]. We analyze the model using the DDD-ITSA feature in the itsa program for Stata [14]. Regression with Newey–West standard errors is specified, with a single lag autoregressive structure (AR1). The Stata post-estimation command lincom is used to compute Wald estimates for linear combinations of coefficients. The exact code used in this example is provided in the Appendix.
3 Results
Table 5 presents the regression results for the DDD-ITSA model. First, we review how well the groups are balanced on baseline level and trend. $\beta_{4}$ and $\beta_{5}$ indicate that California and the primary control states are balanced on baseline level and trend. Linear combinations indicate that California and Colorado are also balanced. Finally, $\beta_{8}$ and $\beta_{9}$ indicate that the two control groups are balanced.
Next, we evaluate treatment effects. $\beta_{7}$ indicates that California experienced a statistically significant annual net reduction of -2.07 per-capita cigarette sales compared to the primary control states. Similarly, the linear combination $(\beta_{7}-\beta_{11})$ indicates a statistically significant annual net decrease relative to the secondary control. The $\beta_{11}$ coefficient indicates that the difference-in-differences between the two control groups was non-statistically significant.
Table 5: Regression results for the DDD-ITSA model
Coefficient Estimate Std Err Z P 95% LCL 95% UCL
$\beta_{0}$ $126.40$ $4.58$ $27.57$ $0.001$ $117.41$ $135.38$
$\beta_{1}$ 1.43-1.43 $0.43$ 3.35-3.35 $0.001$ 2.27-2.27 0.59-0.59
$\beta_{2}$ 11.43-11.43 $4.46$ 2.56-2.56 $0.010$ 20.17-20.17 2.68-2.68
$\beta_{3}$ $0.58$ $0.61$ $0.95$ $0.345$ 0.62-0.62 $1.78$
$\beta_{4}$ $5.83$ $6.20$ $0.94$ $0.347$ 6.33-6.33 $17.98$
$\beta_{5}$ 0.35-0.35 $0.57$ 0.61-0.61 $0.539$ 1.47-1.47 $0.77$
$\beta_{6}$ 8.63-8.63 $6.44$ 1.34-1.34 $0.180$ 21.25-21.25 $3.98$
$\beta_{7}$ 2.07-2.07 $0.75$ 2.77-2.77 $0.006$ 3.54-3.54 0.61-0.61
$\beta_{8}$ $12.29$ $6.84$ $1.80$ $0.072$ 1.11-1.11 $25.70$
$\beta_{9}$ 0.10-0.10 $0.69$ 0.15-0.15 $0.880$ 1.46-1.46 $1.25$
$\beta_{10}$ 6.48-6.48 $7.75$ 0.84-0.84 $0.403$ 21.67-21.67 $8.70$
$\beta_{11}$ 0.31-0.31 $0.88$ 0.35-0.35 $0.723$ 2.03-2.03 $1.41$
  • Note: Standard errors are Newey–West AR(1) adjusted.
Figure 2: Graphic display of the DDD-ITSA outcomes produced by the itsa package in Stata.
4 Discussion
This paper formalizes the triple-difference interrupted time series (DDD-ITSA) design, an extension of the multiple-group ITSA framework that incorporates an additional control group to strengthen causal identification in health policy and intervention research. The DDD-ITSA approach combines the structural time-series logic of ITSA with the econometric rigor of difference-in-differences-in-differences estimators, offering researchers a powerful tool for estimating causal effects when standard two-group comparisons may remain confounded by unmeasured time-varying factors. The key contributions of this paper are the explicit articulation of the DDD-ITSA model, including clear parameter interpretation, and the worked example to illustrate the practical application of the DDD-ITSA design.
Researchers implementing DDD-ITSA face several issues with the design that require careful consideration. First, the choice of primary and secondary control groups is substantively important and should be guided by content expertise and empirical balance assessments. The primary control should be as similar as possible to the treatment unit on observed characteristics and pre-treatment trends [14, 18], while the secondary control provides an additional layer of adjustment, ideally one that shares the primary control’s susceptibility to common shocks but differs in ways that help isolate treatment effects.
Second, when the pool of potential controls is sufficiently large, increasing the number of units assigned to each control group should be considered. Linden [19] found that adding more control units in the MG-ITSA design reduced standard errors and increased power. By extension, incorporating additional comparison groups in a DDD-ITSA would similarly enhance precision by providing more stable estimates of the counterfactual trends across multiple dimensions.
Third, researchers should review autocorrelation and partial autocorrelation functions [3] and formally test for autocorrelation using appropriate diagnostics (e.g., the actest command in Stata [2]) and specify error structures accordingly, whether through Newey–West standard errors [20], Prais–Winsten regression [22], Cochrane–Orcutt regression [5], instrumental variables [11], or other time-series estimators [9, 6].
Fourth, researchers should endeavor to collect as long of time series data as possible. Linden [19] found that in the MG-ITSA context, a longer study period provides the necessary statistical resolution to detect trend-based effects and mitigates the bias introduced by autocorrelation, which can otherwise severely undermine power. Researchers should therefore design DDD-ITSA studies with the expectation that detecting a meaningful trend change will demand a longer time horizon than simpler designs might require.
Finally, Linden [19] found that introducing treatment at the midpoint of the time series maximizes power for identifying trend changes in the MG-ITSA. This would apply equally to DDD-ITSA designs, as balanced pre- and post-treatment periods would minimize standard errors across all interaction terms.
The DDD-ITSA design is particularly well-suited to several common scenarios in health services and policy research. First, when researchers suspect that two-group comparisons may be confounded by concurrent policy changes, economic shocks, or secular trends that differentially affect subgroups, the addition of a second control group can difference out these common sources of bias. Second, DDD-ITSA is valuable when treatment effects are expected to vary across an additional dimension such as demographic subgroups, payer types, or clinical conditions. Third, the design is appropriate when researchers have access to multiple potential control groups but are uncertain which provides the most valid counterfactual.
5 Conclusion
The DDD-ITSA design offers a valuable addition to the health policy researcher’s toolkit, extending the logic of multiple-group interrupted time series to address remaining sources of confounding through the inclusion of a second control group. By differencing across time, groups, and subgroups, DDD-ITSA enhances the credibility of causal claims in observational interrupted time-series settings when suitable control series are available.
Supplementary Information
The Appendix contains Stata code for replicating the Example.
Authors’ contributions
AL conceived the study and its design, conducted all analyses, wrote the manuscript and takes public responsibility for its content.
Funding
There was no funding associated with this work.
Declarations
Ethics approval and consent to participate: Not applicable.
Consent for publication: Not applicable.
Competing interests
The author declares no competing interests.
Appendix
// Download the most recent version of itsa
ssc install itsa, replace
// Example
use "cigsales.dta", replace
// Declare data as time series
tsset state year
* estimate ITSA model in which California is specified as the
* treatment unit (#3 in the data), the treatment period is set
* to 1989, the lag is set to 1 (based on actest), the primary
* control IDs are 8 and 19 (for Idaho and Montana) and the secondary
* control is Colorado (4). We specify that the output should
* include post-treatment trends for all groups, and request that
* a figure be produced
itsa cigsale, treatid(3) trperiod(1989) lag(1) replace ///
contid(8 19) contid2(4) posttrend figure
* baseline level treatment vs control 1 (beta4)
lincom  _b[_z1]
* baseline trend treatment vs control 1 (beta5)
lincom _b[_z1_t]
* baseline level control 2 vs control 1 (beta8)
lincom  _b[_z2]
* baseline trend control 2 vs control 1 (beta9)
lincom _b[_z2_t]
* baseline level treatment vs control 2 (beta4 - beta8)
lincom  _b[_z1] - _b[_z2]
* baseline trend treatment vs control 2 (beta5 - beta9)
lincom _b[_z1_t] - _b[_z2_t]
* DD trends treatment vs control 1 (beta7)
lincom _b[_z1_x_t1989]
* DD trends treatment vs control 2 (beta7 - beta11)
lincom _b[_z1_x_t1989] - _b[_z2_x_t1989]
* DD trends control 1 vs control 2 (beta11)
lincom _b[_z2_x_t1989]
* DDD-ITSA (beta7 - beta11)
lincom _b[_z1_x_t1989] - _b[_z2_x_t1989]